OpenStudy (rainbow_rocks03):

Sole for X. (8 2/5+7 2/3)+4 1/3=x+13

OpenStudy (rainbow_rocks03):

A.6 2/5 B.7 2/5 C.8 2/5 D.16 3/5

OpenStudy (sdfgsdfgs):

Subtract 13 from the left n right hand side of the eqn, what will u get?

OpenStudy (rainbow_rocks03):

IDK will you draw it out for me?

OpenStudy (sdfgsdfgs):

(8 2/5+7 2/3)+4 1/3=x+13 subtract 13 will give u (8 2/5+7 2/3)+4 1/3 - 13 = x + 13 - 13 right?

OpenStudy (rainbow_rocks03):

?

OpenStudy (sdfgsdfgs):

x = (8 2/5+7 2/3) + 4 1/3 - 13 solve for x

OpenStudy (dmndlife24):

Let's simplify what's in parenthesis first... (8 2/5 + 7 2/3) + 4 1/3 = x + 13 241/15 + 4 1/3 = x + 13 Subtract 13 from both sides to get like terms... So we now have 241/15 + 4 1/3 - 13 = x Simplify the left side to get 7 2/5 = x So the answer is B. 7 2/5